# Geometric Embeddability of Complexes is $\exists\mathbb{R}$-complete

### 05 Aug 2021 v1

@@ 2.3 Dimension Reduction
In order to show hardness for all remaining cases of Theorem 1, we establish the following dimension reduction. For interesting dimension reduction in the context of PL-embeddings, we refer to [45, 52, 23, 44].
@@
The idea is to add two apices to a $k$-complex $C$ in order to obtain a $(k+1)$-complex $C^{+}$. We will then argue that $C$ has a geometric embedding in $\mathbb{R}^{d}$ if and only if $C^{+}$ has a geometric embedding in $\mathbb{R}^{d+1}$. More formally, for a complex $C=(V,E)$ and a disjoint vertex set $U$, CUC*U denotes the join complex $(V\cup U,E^{\prime})$ where $E^{\prime}:=\{e\cup u|e\in E,u\in U\}$. The following claim immediately implies Lemma 6.
@@
For the purpose of a contradiction, suppose that $f$ is not injective. Then there exist two distinct points $p=(p_{1},\ldots,p_{d+1})$ and $q=(q_{1},\ldots,q_{d+1})$ with $p,q\in\overline{\varphi}(C)$ such that $(p_{1},\ldots,p_{d})=(q_{1},\ldots,q_{d})$ and $p_{d+1}\neq q_{d+1}$. Without loss of generality, we may assume that $p_{d+1}>q_{d+1}$. Consider the plane $P$ spanned by $\varphi_{a},\varphi_{b},p$. Note that $q\in P$. For an illustration, see Figure 6.
Figure 6: The geometric embedding of $\varphi$ of $C^{+}$, gives a monotone embedding of $C$, otherwise we can find an intersection.
Let us denote with $e_{p}\in E$ and $e_{q}\in E$ any choice of hyperedges such that $p\in\overline{\varphi}(e_{p})$ and $q\in\overline{\varphi}(e_{q})$. Consider the two open segments ${\textrm{seg}}^{\circ}(\varphi_{a},q)\in\overline{\varphi}(e_{q}\cup a)$ and ${\textrm{seg}}^{\circ}(\varphi_{b},p)\in\overline{\varphi}(e_{p}\cup b)$. Clearly, these open segments intersect in a point $x$, as illustrated Figure 6. Because $\varphi$ is a geometric embedding, it holds that
@@ 3 Conclusion
We established the computational complexity of $\textsc{GEM}_{k\rightarrow d}$ for all $d\geq 3$ and $k\in\{d-1,d\}$. In particular, we showed that for these values it is complete for $\exists\mathbb{R}$ to distinguish PL-embeddable $k$-complexes in $\mathbb{R}^{d}$ from geometrically embeddable ones. Arguable, $\textsc{GEM}_{2\rightarrow 3}$ is the most interesting case.
Investigating the computational complexity for the remaining open entries in Table 2 remains for future work. We strengthen the conjecture of Skopenkov [54] as follows.
@@ Acknowledgements.
We thank Arkadiy Skopenkov for his kind and swift help with acquiring literature. Tillmann Miltzow is generously supported by the Netherlands Organisation for Scientific Research (NWO) under project no. 016.Veni.192.250. Linda Kleist was supported by a postdoc fellowship of the German Academic Exchange Service (DAAD).

### 05 Nov 2021 v2

@@ 2.3 Dimension Reduction
In order to show hardness for all remaining cases of Theorem 1, we establish the following dimension reduction. For dimension reductions in the context of PL embeddings, we refer to [undefat, undefaab, undefw, undefas].
@@
The idea is to add two apices to a $k$-complex $C$ in order to obtain a $(k+1)$-complex $C^{+}$. We will then argue that $C$ has a geometric embedding in $\mathbb{R}^{d}$ if and only if $C^{+}$ has a geometric embedding in $\mathbb{R}^{d+1}$. More formally, for a complex $C=(V,E)$ and a disjoint vertex set $U$, CUC*U denotes the join complex $(V\cup U,E^{\prime})$ where $E^{\prime}:=\{e\cup u\mid e\in E,u\in U\}$. The following claim immediately implies Lemma 6.
@@
For the purpose of a contradiction, suppose that $f$ is not injective. Then there exist two distinct points $p=(p_{1},\ldots,p_{d+1})$ and $q=(q_{1},\ldots,q_{d+1})$ with $p,q\in\overline{\varphi}(C)$ such that $(p_{1},\ldots,p_{d})=(q_{1},\ldots,q_{d})$ and $p_{d+1}\neq q_{d+1}$. Without loss of generality, we may assume that $p_{d+1}>q_{d+1}$. Consider the plane $P$ spanned by $\varphi_{a},\varphi_{b},p$. Note that $q\in P$. For an illustration, see Figure 18(a).
Figure 18: Illustration for the proof of Claim 13. The geometric embedding $\varphi$ of $C^{+}$, gives a monotone embedding of $C$, otherwise we can find an intersection.
Let us denote with $e_{p}\in E$ and $e_{q}\in E$ any choice of hyperedges such that $p\in\overline{\varphi}(e_{p})$ and $q\in\overline{\varphi}(e_{q})$. Consider the two open segments ${\textrm{seg}}^{\circ}(\varphi_{a},q)\in\overline{\varphi}(e_{q}\cup a)$ and ${\textrm{seg}}^{\circ}(\varphi_{b},p)\in\overline{\varphi}(e_{p}\cup b)$. Clearly, these open segments intersect in a point $x$, as illustrated Figure 18(a). Because $\varphi$ is a geometric embedding, it holds that
@@ 3 Conclusion
We established the computational complexity of $\textsc{GEM}_{k\rightarrow d}$ for all $d\geq 3$ and $k\in\{d-1,d\}$. In particular, we showed that for these values it is complete for $\exists\mathbb{R}$ to distinguish PL embeddable $k$-complexes in $\mathbb{R}^{d}$ from geometrically embeddable ones. Arguably, $\textsc{GEM}_{2\rightarrow 3}$ is the most interesting case.
Investigating the computational complexity for the remaining open entries in Table 2 remains for future work. We strengthen the conjecture of Skopenkov [undefaad] as follows.
@@ Acknowledgements.
We thank Arkadiy Skopenkov for his kind and swift help with acquiring literature. We thank Martin Tancer for pointing out a mistake in a previous version of this manuscript. Mikkel Abrahamsen is part of Basic Algorithms Research Copenhagen (BARC), generously supported by the VILLUM Foundation grant 16582. Linda Kleist is generously supported by a postdoc fellowship of the German Academic Exchange Service (DAAD). Tillmann Miltzow is generously supported by the Netherlands Organisation for Scientific Research (NWO) under project no. 016.Veni.192.250.
Table 3: A glossary for notions used in the proof of Theorem 4.
symbol meaning
$L=\{\ell_{1},\ldots,\ell_{n}\}$ pseudoline arrangement
$n$ number of pseudolines in $L$, Figure 4(a)
$\ell_{0}$ additional pseudoline
$^{*}:=L\cup\{\ell_{0}\} LL$ together with $\ell_{0}$
$\triangle$ the helper triangle, all intersections of $L^{*}$ are contained inside
$C=(V,E)$ the simplicial complex that we construct
$T^{+}_{i}$ tunnel around pseudoline $\ell_{i}$, Figure 5(a)
$T_{i}$ tunnel $T^{+}_{i}$ without first and last two sections
section part of tunnel, Figure 5(b)
tunnel bottom the part of tunnel $T^{+}_{i}$ that lies in the plane $z=i$ in step 1
tunnel roof the part of tunnel which is not in the bottom, see Figure 5(c)
left/right roof path tunnel paths shared by roof and bottom.
central roof path tunnel path that is disjoint from left/right roof path
$e_{i}=(s_{i},t_{i})$ special edge of $C$ that is meant to represent $\ell_{i}$
$u$ apex (taking the role of an observer high above)
$H$ graph in the third step of the construction, see black graph in Figure 7
$G^{+}$ an essentially $3$-connected planar graph induced by the neighborhood of $u$
$f_{0}$ outer face of $G^{+}$
$G$ subgraph of $G^{+}$ that is inside the helper triangle
$\overline{C}$ the complex $C$ consists of two copies of $\overline{C}$ with the apex identified
$\varphi$ geometric embedding of $C$
tent Figure 9(b)
tunnel loop $o$ triangle that is shared by two sections
rafter two roof edges of a tunnel loop
wedge $W_{i}$ defined by the apex $u$ and $e_{i}$, Figure 10
plane $P_{i}$ defined by the apex $u$ and $e_{i}$
$p>e_{i}$ segment $pu$ does not cross $e_{i}$
$p<e_{i}$ segment $pu$ does cross $e_{i}$
$S$ sphere around apex $u$
$D$, $D^{+}$ projection of $\varphi(G^{+})$ onto $S$ yielding a drawing of $G$ and $G^{+}$
$a_{i}$ projection of $\varphi(e_{i})$ onto $S$
$\mathcal{A}$ arc arrangement of all $a_{i}$ Figure 11(b)